A good vacuum produced in common laboratory apparatus corresponds to 10^-6mm pressure at 25℃. Calculate no. Of molecules per cc at this pressure and temperature.(Ans:-3.2×10^10)

Grade:11

4 Answers

Vikas TU
7 years ago
From PV =nRT
1.3 x 10^-9 * 10^-3 = n * 0.0821 * 300
n = 0.0527 x 10^-12  
no. of molecules = n x Na => 0.0527 x 10^-12  x 6.023 x 10^23 = > 0.317 x 10^11 
or => 3.17 x 10^10 
Vaibhavi
7 years ago
How you got the volume? Plz explain the answer in descriptive manner.
Gaurav
7 years ago
1 ml divided by 1000 ml/ liter divided by 24.1 liters per mole multiplied by 6.023 molecules per mole divided by 760 mm Hg multiplied by 1 x 10^-6 mm Hg = molecule per ml at the reported temperature and pressure reported as one significant figure..
Ankita Mishra
6 years ago
10^-6mm=10^-6/760atm
volume=1l=1000c.c
PV=nRT
10^-6/760atm*1000=n.*0821*298(25^0 c=298k)
n=10^-6/760atm*1000 / 0821*298
n=0.0538*10^-6 mole per c.c
no .of moleclue per c.c=0.0538*10^-6 * 6.022 * 10^23
                                    =3.2*10^16